OpenStudy (anonymous):

Which of the following equations represents the graph? f(x) = (x - 4)2 - 6 f(x) = (x + 4)2 + 6 f(x) = (x + 4)2 - 6 f(x) = (x - 4)2 + 6

OpenStudy (anonymous):

this is the graph

ganeshie8 (ganeshie8):

start by finding the coordinates of vertex, the turning point

OpenStudy (anonymous):

-6, -4??

ganeshie8 (ganeshie8):

x coordinate always comes first vertex = (h, k) = (-4, -6) ok ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

remember the vertex form of parabola ?

ganeshie8 (ganeshie8):

y = a(x - h)^2 + k plug coordinates of vertex above

OpenStudy (anonymous):

y = a(x-4)^2-6?

ganeshie8 (ganeshie8):

careful, h is negative 4

ganeshie8 (ganeshie8):

it should be y = a(x- (-4))^2-6 which is same as y = a(x + 4)^2 - 6

OpenStudy (anonymous):

right so its like flipping it

ganeshie8 (ganeshie8):

look at the options

OpenStudy (anonymous):

OH thank you so much!

ganeshie8 (ganeshie8):

yw!